# The H2M Monolithic Active Pixel Sensor — characterizing non-uniform in-pixel response in a 65 nm CMOS imaging technology

### v1

@@ Abstract
The high energy physics community recently gained access to the TPSCo ISC (Image Sensor CMOS), which enables a higher in-pixel logic density in monolithic active pixel sensors (MAPS) compared to processes with larger feature sizes. To explore this novel technology, the Hybrid-to-Monolithic (H2M) test chip has been designed and manufactured. The design followed a digital-on-top design workflow and ports a hybrid pixel-detector architecture, with digital pulse processing in each pixel, into a monolithic chip. The chip matrix consists of 64×16 square pixels with a size of , and a total active area of approximately . The chip has been successfully integrated into the Caribou DAQ system. It is fully functional, and the measured threshold dispersion and noise agree with the expectation from front-end simulations. However, a non-uniform in-pixel response related to the size and location of the n-wells in the analog circuitry has been observed in test beam measurements and will be discussed in this contribution. This asymmetry in the pixel response, enhanced by the pixel pitch — larger than in other prototypes — and certain features of the readout circuit, has not been observed in prototypes with smaller pixel pitches in this technology.
@@ 1 Introduction
Monolithic CMOS sensors enable the development of detectors with a low material budget and a low fabrication cost. By using a small collection electrode, these sensors achieve small capacitance, low analog power consumption, and an improved signal-to-noise ratio. The availability of a CMOS imaging process to the high-energy physics community further enhances their potential by allowing a higher density of in-pixel logic in monolithic active pixel sensors (MAPS) compared to processes with larger feature sizes. This process has been previously explored in prototypes such as the APTS [1] and DPTS [2].
The H2M (Hybrid-to-Monolithic) test chip has been developed to demonstrate the capabilities of the CMOS imaging process in tracking applications for future high-energy lepton experiments, while also addressing the challenges of porting a hybrid architecture into a monolithic chip, and testing a compact digital cell library and a complex and fast in-pixel readout circuitry.
@@ 2 The H2M test chip
The H2M sensor is manufactured using a modified TPSCo ISC (Image Sensor CMOS). It comprises a small n-type collection electrode placed on a high-resistivity p-type epitaxial layer, which is grown on a low-resistivity substrate. To enhance charge collection in the sensitive layer with a thickness of , process modifications have been implemented, including a low-dose n-type implant with a gap at the pixel boundaries [3]. The wafers are backside-thinned to physical thickness. A schematic of the sensor layout is illustrated in figure 1.
The chip matrix consists of 6416 square pixels with a size of , resulting in a total active area of approximately . Each pixel includes an analog front end that comprises a charge-sensitive amplifier (CSA) with Krummenacher feedback [4] followed by a discriminator. The design ensures a constant slope of the falling edge of the CSA output, adjustable through the Krummenacher feedback current (ikrum). The global threshold of the discriminator is selected with an 8-bit DAC, and threshold mismatch is compensated using a 4-bit threshold-tuning DAC per pixel. Individual pixels can be masked, and if needed, test pulses of varying amplitudes can be injected before the CSA. The digital logic processes the output of the discriminator. Both the CMOS analog and digital front ends are hosted within the deep p-wells. The analog front-end begins at the collection electrode and extends in the row direction, with the remaining space occupied by the digital readout circuitry. PMOS devices are shielded in n-wells inside the deep p-well.
@@ 3 Performance in test beam
Figure 2(a) shows the efficiency and fake hit rate as a function of the hit detection threshold for three different bias voltages. The bias voltage on the p-well/substrate influences the depletion of the sensor volume. At low bias voltages (), the area surrounding the collection electrode is not fully depleted, which increases the sensor capacitance and, consequently, noise and fake hit rate. Allpix2 simulations, using the results from Technology computer-aided design (TCAD) simulations based on generic profiles [10], are performed to obtain the efficiency as a function of the detection threshold, which is also included in figure 2(a). The difference between the measured and simulated efficiency can be explained by a non-uniformity in the in-pixel response, which will be further discussed below.
Figure 3(a) and 3(b) show the efficiency and ToA as a function of the in-pixel hit position. These maps are obtained by projecting all track intercepts into four pixels. High efficiency and fast charge carrier collection occur near the collection electrode but decrease asymmetrically towards the edges and corners. Particularly, the drop in charge collection speed, resulting in reduced efficiency, is located beneath the n-well of the analog front-end (see figure 1). While the n-wells in the digital readout circuitry are thin and uniformly distributed, the n-well in the analog front-end has a width of about , which affects the electric field at the interface of the p-well and the low-dose n-implant layer, creating local potential wells that slow down charge collection. This does not influence charge sharing, and the cluster size map is therefore symmetric, as shown in figure 3(c).
@@
Figure 4 illustrates the CSA output for slow and fast signals for large and small ikrum. The loss of the CSA output signal height for slow signals occurs due to a mismatch between the charge collection time and the CSA response time. This effect is known as ballistic deficit. While the signal amplitude for the fast signals barely depends on ikrum, slow signals yield a different amplitude. As a result, a low ikrum improves the overall chip efficiency, as demonstrated in figure 2(b). At low hit detection thresholds and larger bias voltages, the in-pixel response remains uniform, achieving an efficiency of at a threshold of with the sensor biased at .
This effect is influenced by the fast analog front-end and the large pitch of H2M [11]. The integration time of a few nanoseconds contributes to the ballistic deficit by reducing the effective amplitude of slower signals compared to faster ones. Additionally, the large pitch of results in significant spacing between the lateral electric field components of the collection electrodes and the gap, which makes it more sensitive to field perturbations due to the n-wells in the p-well region. This makes H2M the only chip in this technology where such non-uniform in-pixel responses have been observed.

### v2

@@ Abstract
The high energy physics community recently gained access to the TPSCo ISC (Image Sensor CMOS), which enables a higher in-pixel logic density in monolithic active pixel sensors (MAPS) compared to processes with larger feature sizes. To explore this novel technology, the Hybrid-to-Monolithic (H2M) test chip has been designed and manufactured. The design follows a digital-on-top design workflow and ports a hybrid pixel-detector architecture, with digital pulse processing in each pixel, into a monolithic chip. The chip matrix consists of 6416 square pixels with a size of , and a total active area of approximately . The chip is operated and read out using the Caribou DAQ system. The measured threshold dispersion and noise agree with the expectation from front-end simulations. However, a non-uniform in-pixel response related to the size and location of the n-wells in the analog circuitry has been observed in test beam measurements and will be discussed in this contribution. This asymmetry in the pixel response, enhanced by the pixel pitch — larger than in other prototypes — and certain features of the readout circuit, has not been observed in prototypes with smaller pixel pitches in this technology.
@@ 1 Introduction
Monolithic CMOS sensors enable the development of detectors with a low material budget and a low fabrication cost compared to hybrid pixels. By employing a small collection electrode, these sensors achieve small input capacitance, low analog power consumption, and an improved signal-to-noise ratio compared to sensors with larger collection electrodes. The availability of a CMOS imaging process to the high-energy physics community further enhances the potential of monolithic active pixel sensors (MAPS), as it allows for a higher density of in-pixel logic compared to processes with larger feature sizes. This process has been previously explored in prototypes such as the APTS [1] and DPTS [2].
The H2M (Hybrid-to-Monolithic) test chip has been developed to demonstrate the capabilities of the CMOS imaging process in tracking applications for future high-energy lepton experiments (such as CLIC or FCCee). The architecture of this monolithic chip is based on a hybrid readout chip (from the Timepix family), with the goal of learning about the challenges and potential associated with the porting process. The development aimed to design and evaluate a set of digital logic building blocks, a so-called digital cell library, with a smaller footprint (reduction of 25% in logic area compared to other available standard cell libraries), as well as complex and fast in-pixel readout circuitry explained below.
@@ 2 The H2M test chip
The H2M sensor is manufactured using a modified TPSCo ISC (Image Sensor CMOS) [3]. It comprises a small n-type collection electrode placed on a high-resistivity p-type epitaxial layer, which is grown on a low-resistivity substrate. To enhance charge collection in the sensitive layer with a thickness of , process modifications have been implemented, including a low-dose n-type implant with a gap at the pixel boundaries [3]. The wafers are backside-thinned to physical thickness. A schematic of the sensor layout is illustrated in figure 1.
The chip matrix consists of 6416 square pixels with a size of , resulting in a total active area of approximately . Each pixel includes an analog front end that comprises a charge-sensitive amplifier (CSA) with Krummenacher feedback [4] followed by a discriminator. The design ensures a constant slope of the falling edge of the CSA output, adjustable through the feedback current (ikrum), where a lower ikrum effectively increases the integration time, enhancing the signal and improving the signal efficiency for a given threshold. The global threshold of the discriminator is selected with an 8-bit DAC, and threshold mismatch is compensated using a 4-bit threshold-tuning DAC per pixel. Individual pixels can be masked, and if needed, test pulses of varying amplitudes can be injected at the input stage of the CSA. The digital logic processes the output of the discriminator. Both the CMOS analog and digital front ends are hosted within the deep p-wells. The analog front-end begins at the collection electrode and extends in the row direction, with the remaining space occupied by the digital readout circuitry. PMOS devices are shielded in n-wells inside the deep p-well.
@@ 3 Performance in test beam
Figure 2(a) shows the efficiency and fake hit rate as a function of the hit detection threshold for three different bias voltages measured in triggered mode with a fixed frame duration of per event. The bias voltage on the p-well/substrate influences the depletion of the sensor volume. At low bias voltages (), the area surrounding the collection electrode is not fully depleted, which increases the sensor capacitance and, consequently, noise and fake hit rate. Allpix2 simulations, using the results from Technology computer-aided design (TCAD) simulations based on generic profiles [10], are performed to obtain the efficiency as a function of the detection threshold, which is also included in figure 2(a). The difference between the measured and simulated efficiency can be explained by a non-uniformity in the in-pixel response, which will be further discussed below.
Figure 3(a) and 3(b) show the efficiency and ToA as a function of the in-pixel hit position. These maps are obtained by projecting all track intercepts into four pixels. High efficiency and fast charge carrier collection occur near the collection electrode but decrease asymmetrically towards the edges and corners. Particularly, the drop in charge collection speed, resulting in reduced efficiency, is located beneath the n-well of the analog front-end (see figure 1). While the n-wells in the digital readout circuitry are thin and uniformly distributed, the n-well in the analog front-end has a width of about , which affects the electric field at the interface of the p-well and the low-dose n-implant layer, creating local potential wells that slow down charge collection along their paths (see figure 1). This does not influence charge sharing, and the cluster size map is therefore symmetric, as shown in figure 3(c).
@@
Figure 4 illustrates the CSA output for slow and fast signals for large and small ikrum. The loss of the CSA output signal height for slow signals occurs due to a mismatch between the charge collection time and the CSA response time. This effect is known as ballistic deficit. While the signal amplitude for the fast signals barely depends on ikrum, slow signals yield a different amplitude. As a result, a low ikrum improves the overall chip efficiency, as demonstrated in figure 2(b). At low hit detection thresholds (< ) and larger bias voltages, the in-pixel response remains uniform, achieving an efficiency of at a threshold of with the sensor biased at .
This effect is enhanced by the fast analog front-end and the large pitch of H2M [11]. The integration time of a few nanoseconds contributes to the ballistic deficit by reducing the amplitude of slower signals compared to faster ones. Additionally, the relatively large pitch of results in significant spacing between the lateral electric field components of the collection electrodes and the gap, which makes it more sensitive to field perturbations due to the n-wells in the p-well region. This makes H2M the only chip in this technology where such non-uniform in-pixel responses have been observed.